40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8424     Entry time: Mon Apr 8 22:43:44 2013
Author: rana 
Type: Update 
Category: PSL 
Subject: MC locking troubles: MC/FSS servo unstable 

The MC seemed to be losing lock recently quite a bit. I noticed that the PC Drive RMS signal was red.

This means that the high frequency drive to the Pockels cell was too high by a factor of 2-3 and sometimes saturating and breaking the lock.

I fiddled with the gains on the FSS screen until this value went down. It looks like there is some kind of high Q oscillation; it takes a couple minutes for the PC Drive RMS to settle to its new position after changing the gains.

The attached trend plot show the last 2 hours. The mean is now back to ~1 V and seems OK. We should really examine the FSS or MC error point spectra with the RF analyzer while exploring this gain space.

Attachment 1: Untitled.png  23 kB  Uploaded Tue Apr 9 00:19:46 2013  | Hide | Hide all
Untitled.png
ELOG V3.1.3-